OpenStudy (anonymous):

Solve for . 22=2U-8 Simplify your answer as much as possible.

OpenStudy (anonymous):

u=15?

OpenStudy (anonymous):

Or 11=U-4?

OpenStudy (anonymous):

22=2U-8 22+8=2U U=22+8/2 U=11+4 U=15

OpenStudy (anonymous):

Alright I will gve u a medal

OpenStudy (anonymous):

22=2U-8 22+8=2U U=22+8/2 U=11+4 U=15 SAME AS 126 HE IS RIGHT!!!!!!!!!!!!!!!!!!!!